OpenStudy (anonymous):

If the cartesian equation of a line are 3 - 5/5 = y + 4/7 = 2z -6/4 Write the vector equation for a line vector r = vector a + lambda(vector b)

ganeshie8 (ganeshie8):

\[\large \dfrac{x-5}{5} = \dfrac{y+4}{7} = \dfrac{2z-6}{4}\]

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

Sorry internet problem, I got it!  @ganeshie8